# Differentially Private Bilevel Optimization

### v2

@@ 5 Mini-batch algorithm for DP bilevel ERM
In this section, we consider once again the ERM bilevel problem, (BO/ERM), and provide Algorithm 3, which is a mini-batch variant of the ERM algorithm discussed in the previous section. Given a mini-batch $B\subseteq S=\{\xi_{1},\dots,\xi_{n}\}$ and a function $h:\mathbb{R}^{d}\times\Xi\to\mathbb{R}$, we let $\nabla h(\mathbf{z};B)=\frac{1}{|B|}\sum_{\xi_{i}\in B}\nabla h(\mathbf{z};\xi_{i})$ denote the mini-batch gradient. We prove the following result:
@@
Assume 2.5 and 2.6 hold, and that $\alpha\leq\ell\kappa^{3}\min\{\frac{1}{2\kappa},\frac{L_{0}^{g}}{L_{0}^{f}},\frac{L_{1}^{g}}{L_{1}^{f}},\frac{\Delta_{F}}{\ell\kappa}\}$. Then running Algorithm 3 with parameters assigned as in Theorem 4.1 and any batch sizes $b_{\mathrm{in}},b_{\mathrm{out}}\in[n]$ for sampling gradients, satisfies $(\epsilon,\delta)$-DP and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
@@
where $K_{1}=\mathcal{O}(\Delta_{F}^{1/4}\ell^{3/4}\kappa^{5/4}),~{}K_{2}=\mathcal{O}(\Delta_{F}^{1/6}\ell^{1/2}\kappa^{11/6}),~{}K_{3}=\mathcal{O}(\ell\kappa)$.
@@
5:     $\mathbf{x}_{t+1}=\arg\min_{\mathbf{u}\in\mathcal{X}}\left\{\left\langle\widetilde{\mathbf{g}}_{t},\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}_{t}\right\|^{2}\right\}$triangleright If $\mathcal{X}=\mathbb{R}^{d_{x}}$, then $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\widetilde{\mathbf{g}}_{t}$
@@
7:$t_{\mathrm{out}}=\arg\min_{t\in\{0,\dots,T-1\}}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|$.
@@
Remark 5.2 (Outer batch size dependence).
Algorithm 3 ensures privacy for any batch sizes, yet notably, the guaranteed gradient norm bound does not go to zero (as $n\to\infty$) for constant outer-batch size. The same phenomenon also holds for for “single”-level constrained nonconvex optimization, as noted by Ghadimi et al. (2016) (specifically, see Corollary 4 and related discussion therein). Accordingly, the inner-batch size $b_{\mathrm{in}}$ can be set whatsoever, while setting $b_{\mathrm{out}}=\mathcal{O}(\max\{(\epsilon n/\sqrt{d_{x}})^{1/2},(\epsilon n/\sqrt{d_{y}})^{1/3}\})\ll n$ recovers the full-batch rate (in general, $\lim_{n\to\infty}\|\mathcal{G}(\mathbf{x}_{\mathrm{out}})\|=0$ whenever $b_{\mathrm{out}}=\omega_{n}(1)$ grows with the sample). It is interesting to note that the $1/b_{\mathrm{out}}$ term shows up in the analysis only as an upper bound on the sub-Gaussian norm of the mini-batch gradient estimator. Thus, in applications for which some (possibly small) batch size results in reasonably accurate gradients, the result above holds with the outer mini-batch gradient’s standard deviation replacing $1/b_{\mathrm{out}}$, which is to be expected anyhow for high probability guarantees.
The difference between Algorithm 3 and Algorithm 1, is that both the inner and outer loops sample mini-batch gradients. The inner loop guarantee is the same regardless of the inner-batch size $b_{\mathrm{in}}$, since for strongly-convex objectives it is possible to prove the same convergence rate guarantee for DP optimization in any case (as further discussed in Appendix A). As to the outer loop (Line 5), we apply standard concentration bounds to argue about the quality of the gradient estimates — hence the additive $1/b_{\mathrm{out}}$ factor — and rely on our analysis of the outer loop with inexact gradients (which are now even less exact due to sampling stochasticity). We remark that compared to the well-known analysis of Ghadimi et al. (2016) for constrained nonconvex optimization, we derive high probability bounds without requiring several re-runs of the algorithm. We further remark that mini-batch sampling with replacement is analyzed for simplicity, though the same guarantees (up to constants) can be derived for sampling without replacement, at the cost of a more involved analysis.
@@ 6 Generalizing from ERM to population loss
Under Assumptions 2.5 and 2.6, if the preconditions of Theorem 4.1 hold, then Algorithm 1 is $(\epsilon,\delta)$-DP, and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
$\left\|\mathcal{G}_{F_{\mathcal{P}},\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha=\widetilde{\mathcal{O}}\bigg{(}K_{1}\Big{(}\frac{\sqrt{d_{x}}}{\epsilon n}\Big{)}^{1/2}+K_{2}\Big{(}\frac{\sqrt{d_{y}}}{\epsilon n}\Big{)}^{1/3}+K_{3}\Big{(}\frac{d_{x}}{n}\Big{)}^{1/2}\bigg{)}~{},$
where $K_{1}=\mathcal{O}(\Delta_{F}^{1/4}\ell^{1/4}\kappa^{5/4}),~{}K_{2}=\mathcal{O}(\Delta_{F}^{1/6}\ell^{1/2}\kappa^{11/6}),~{}K_{3}=\mathcal{O}(\ell\kappa)$. Similarly, if the preconditions of Theorem 5.1 hold, then for any batch sizes $b_{\mathrm{in}},b_{\mathrm{out}}\in[n]$, Algorithm 3 is $(\epsilon,\delta)$-DP, and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
$\left\|\mathcal{G}_{F_{\mathcal{P}},\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha=\widetilde{\mathcal{O}}\bigg{(}K_{1}\Big{(}\frac{\sqrt{d_{x}}}{\epsilon n}\Big{)}^{1/2}+K_{2}\Big{(}\frac{\sqrt{d_{y}}}{\epsilon n}\Big{)}^{1/3}+K_{3}\cdot\frac{1}{b_{\mathrm{out}}}+K_{3}\Big{(}\frac{d_{x}}{n}\Big{)}^{1/2}\bigg{)}~{}.$
The proof of Theorem 6.1 relies on arguing that the hyperobjective is Lipschitz, and applying a uniform convergence bound for bounded gradients, which we further implies uniform convergence of projected gradients by Lemma 2.4.
@@ 7 Application: private regularization hyperparameter tuning
A second approach to hyperparameter tuning is based on differentiable programming: instead of predetermining a finite set of possible values for $\omega$, we can solve Eq. (4) using bilevel optimization methods, namely “differentiate through $\omega$”, and update it accordingly on the fly. This approach was pioneered by Bengio (2000), and was further developed by multiple works (Maclaurin et al., 2015; Franceschi et al., 2017, 2018; Lorraine et al., 2020; Engstrom et al., 2025). However, we are not aware of a private variant of this methodology, and our goal here is to provide one.
@@
$\omega_{t+1}$ $=\arg\min_{u\geq 0}\left\{u\cdot\left[\nabla_{\omega}f(\omega_{t},\theta_{t}^{\lambda})+\lambda(\nabla_{\omega}g(\omega_{t},\theta_{t}^{\lambda})-\nabla_{\omega}g(\omega_{t},\theta_{t}))+\mathcal{N}(0,\sigma^{2})\right]+\tfrac{1}{2\eta}(u-\omega_{t})^{2}\right\}$
$=\arg\min_{u\geq 0}\left\{u\cdot\left[0+\lambda(R(\theta_{t}^{\lambda})-R(\theta_{t}))+\mathcal{N}(0,\sigma^{2})\right]+\tfrac{1}{2\eta}(u-\omega_{t})^{2}\right\}$
@@ 8 Proofs
Note that the two sub-problems solved by Line 3 of Algorithm 1 are strongly-convex and admit Lipschitz components over $\mathcal{Y};~{}g(\mathbf{x},\cdot)$ by assumption, and $f+\lambda g$ by combining this with the smoothness/Lipschitzness of $f$, as follows:
@@
Although the rate in Theorem 8.2 appears in prior works such as (Bassily et al., 2014; Feldman et al., 2020), it is typically manifested through a bound in expectation (and in terms of function value) as opposed to with high probability, required for our purpose. We therefore, for the sake of completeness, provide a self-contained proof of Theorem 8.2 in Appendix A.
@@
If $\lambda\geq\max\{\frac{2L_{1}^{g}}{\mu_{g}},\frac{L_{0}^{f}}{L_{0}^{g}}\}$, then $\widetilde{\mathbf{y}}_{t}$ and $\widetilde{\mathbf{y}}_{t}^{\lambda}$ (as appear in Line 3 of Algorithm 1) satisfy with probability at least $1-\gamma:$
@@
where the second bound holds for $\lambda\geq\frac{L_{1}^{f}}{L_{1}^{g}}$ ∎
@@
Noting that $\nabla^{2}_{xy}\mathcal{L}_{\lambda}\preceq 2\lambda L_{1}^{g}$ and $\nabla^{2}_{yy}\mathcal{L}_{\lambda}\succeq\lambda\mu_{g}/2$ everywhere, hence $[\nabla^{2}_{yy}\mathcal{L}_{\lambda}]^{-1}\preceq 2/\lambda\mu_{g}$ we get that
@@
As $\nu_{0},\dots,\nu_{T-1}\overset{iid}{\sim}\mathcal{N}(0,\sigma^{2}I)$, a standard Gaussian norm bound (cf. Vershynin 2018, Theorem 3.1.1) ensures that with probability at least $1-\gamma$, for all $t\in\{0,1,\dots,T-1\}:~{}\left\|\nu_{t}\right\|^{2}\lesssim d\sigma^{2}\log(T/\gamma)\lesssim\frac{\alpha^{2}}{64}$. We therefore condition the rest of the proof on the highly probable event under which this uniform norm bound indeed holds. We continue by introducing some notation. We denote ~$\widetilde{\nabla}_{t}=\widetilde{\nabla}{h}(\mathbf{x}_{t})+\nu_{t}$, and $\delta_{t}:=\widetilde{\nabla}_{t}-\nabla h(\mathbf{x}_{t})$. We further denote
@@
It holds that ~$\langle\widetilde{\nabla}_{t},\rho_{t}\rangle\geq\left\|\rho_{t}\right\|^{2}$.
@@
~$\left\langle\widetilde{\nabla}_{t}+\frac{1}{\eta}(\mathbf{x}_{t+1}-\mathbf{x}_{t}),\mathbf{u}-\mathbf{x}_{t+1}\right\rangle\geq 0~{}.$
@@
With the lemmas above in hand, we are now ready to prove Proposition 4.6. Note that by construction, the algorithm returns the index $t$ with minimal $\left\|\rho_{t}\right\|$. Further note that $\left\|\rho_{t}-\mathcal{G}_{t}\right\|\leq\left\|\delta_{t}\right\|$ by Lemma 2.4, thus
@@
where the second inequality is due to Lemma 2.4, the third due to Lemma 4.3.b, and the last by our assignment of $\lambda$. It therefore remains to bound $\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|$.
@@
$\beta=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}\right)=\widetilde{\mathcal{O}}\left(\frac{\ell^{3/2}\kappa^{11/2}\Delta_{F}^{1/2}\sqrt{d_{y}}}{\alpha^{2}\epsilon n}\right)$ (12)
@@
If $\lambda\geq\max\{\frac{2L_{1}^{g}}{\mu_{g}},\frac{L_{0}^{f}}{L_{0}^{g}},\frac{L^{f}_{1}}{L^{g}_{1}}\}$, then there is $\beta_{b}=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}+\frac{\ell\kappa}{b}\right)$ such that with probability at least $21-\gamma/$ 2, $\mathbf{g}_{t}^{B}:=\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};B_{t})+\lambda\left(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};B_{t})-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t};B_{t})\right)$ satisfies for all $t\in\{0,\dots,T-1\}:~{}\|\nabla\mathcal{L}^{*}_{\lambda}(\mathbf{x}_{t})-\mathbf{g}^{B}_{t}\|\leq\beta_{b}$.
@@
following the same proof as Lemma 4.4 in Section 8.1, by replacing Theorem 8.2 by the mini-batch Theorem A.1 (whose guarantee holds regardless of the inner batch size) .
To bound the second summand, note that $\|\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};\xi)+\lambda(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};\xi)-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t};\xi))\|\leq M=\mathcal{O}(\ell\kappa)$ for every $\xi\in\Xi$, by Lemma 4.5. Hence, $\mathbf{g}^{B}_{t}$ is the average of $b$ independent vectors bounded by $M$, all with the same mean, and therefore a standard concentration bound (cf. Jin et al. 2019) ensures that $\|\mathbf{g}_{t}^{B}-\mathbb{E}[\mathbf{g}_{t}^{B}]\|=\widetilde{\mathcal{O}}(M/b)$ with probability at least $41-\gamma/$ 4, which completes the proof. ∎
@@
We start by proving the privacy guarantee. Since $\mathcal{L}^{*}_{\lambda,i}$ is $\mathcal{O}(\ell\kappa)$-Lipschitz by Lemma 4.5, the sensitivity of $\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};B_{t})+\lambda\left(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};B_{t})-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t};B_{t})\right)$ is at most $\mathcal{O}(\ell\kappa)$. Accordingly, the “unamplified” Gaussian mechanism (Lemma 2.2) ensures $(\tilde{\epsilon},\tilde{\delta})$-DP for $\tilde{\epsilon}=\widetilde{\Theta}\left(\frac{\ell\kappa}{b\sigma}\right)$, and hence is amplified (Lemma 2.3) to $(\epsilon_{0},\delta_{0})$-DP for $\epsilon_{0}=\widetilde{\Theta}\left(\frac{\ell\kappa}{b\sigma}\cdot\frac{b}{n}\right)=\frac{\epsilon}{\sqrt{18T}}$, the last holding for sufficiently large $\sigma^{2}=\widetilde{\Theta}\left(\frac{\ell^{2}\kappa^{2}T}{\epsilon^{2}n^{2}}\right)$, and for $\delta_{0}=\frac{\delta}{3(T+1)}$. Therefore, basic composition shows that each iteration of the algorithm is $(3\cdot\frac{\epsilon}{\sqrt{18T}},3\frac{\delta}{3(T+1)})=(\frac{\epsilon}{\sqrt{2T}},\frac{\delta}{T+1})$-DP. Since $\epsilon/\sqrt{T}\ll 1$ under our parameter assignment, advanced composition over the $T$ iterations yields the $(\epsilon,\delta)$-DP guarantee.
@@
where the second inequality is due to Lemma 2.4, the third due to Lemma 4.3, and the last by our assignment of $\lambda$. It therefore remains to bound $\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|$.
@@
$\beta_{b}=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}+\frac{\ell\kappa}{b}\right)=\widetilde{\mathcal{O}}\left(\frac{\ell^{3/2}\kappa^{11/2}\Delta_{F}^{1/2}\sqrt{d_{y}}}{\alpha^{2}\epsilon n}+\frac{\ell\kappa}{b}\right)~{},$ (16)
@@
We first need a simple lemma that immediately follows from our assumptions, and Eq. (1):
@@
Under Assumptions 2.5 and 2.6, $F(\cdot\,;\xi)$ is $G$-Lipschitz, for $G=\mathcal{O}(\ell\kappa)$.
Accordingly, the main tool that will allow us to obtain generalization guarantees, is the following uniform convergence result in terms of gradients:
Lemma 8.10 (Mei et al., 2018, Theorem 1).

### v3

@@ 5 Mini-batch algorithm for DP bilevel ERM
In this section, we consider once again the bilevel ERM problem (BO/ERM), and provide Algorithm 3, which is a mini-batch variant of the previously discussed bilevel ERM algorithm. Given a mini-batch $B\subseteq S=\{\xi_{1},\dots,\xi_{n}\}$ and a function $h:\mathbb{R}^{d}\times\Xi\to\mathbb{R}$, we let $\nabla h(\mathbf{z};B)=\frac{1}{|B|}\sum_{\xi_{i}\in B}\nabla h(\mathbf{z};\xi_{i})$ denote the mini-batch gradient. We prove the following result:
@@
Assume 2.1 and 2.2 hold, and that $\alpha\leq\ell\kappa^{3}\min\{\frac{1}{2\kappa},\frac{L_{0}^{g}}{L_{0}^{f}},\frac{L_{1}^{g}}{L_{1}^{f}},\frac{\Delta_{F}}{\ell\kappa}\}$. Then running Algorithm 3 with parameters assigned as in Theorem 4.1 and any batch sizes $b_{\mathrm{in}},b_{\mathrm{out}}\in[n]$ for sampling gradients, satisfies $(\epsilon,\delta)$-DP and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
@@
where $K_{1}=\mathcal{O}(\Delta_{F}^{1/4}\ell^{3/4}\kappa^{5/4}),~K_{2}=\mathcal{O}(\Delta_{F}^{1/6}\ell^{5/6}\kappa^{11/6}),~K_{3}=\mathcal{O}(\ell\kappa)$.
Remark 5.2 (Outer batch size dependence).
Algorithm 3 ensures privacy for any batch sizes, yet notably, the guaranteed gradient norm bound does not go to zero (as $n\to\infty$) for constant outer-batch size. The same phenomenon also holds for for “single”-level constrained nonconvex optimization, as noted by Ghadimi et al. (2016) (specifically, see Corollary 4 and related discussion therein). Accordingly, the inner-batch size $b_{\mathrm{in}}$ can be set whatsoever, while setting $b_{\mathrm{out}}=\mathcal{O}(\max\{(\epsilon n/\sqrt{d_{x}})^{1/2},(\epsilon n/\sqrt{d_{y}})^{1/3}\})\ll n$ recovers the full-batch rate (in general, $\|\mathcal{G}(\mathbf{x}_{\mathrm{out}})\|\overset{n\to\infty}{\longrightarrow}0$ whenever $b_{\mathrm{out}}\overset{n\to\infty}{\longrightarrow}\infty$). It is interesting to note that the $1/b_{\mathrm{out}}$ term shows up in the analysis only as an upper bound on the sub-Gaussian norm of the mini-batch gradient estimator. Thus, in applications for which some (possibly small) batch size results in reasonably accurate gradients, the result above holds with the outer mini-batch gradient’s standard deviation replacing $1/b_{\mathrm{out}}$, which is to be expected anyhow for high probability guarantees.
@@
5:  $\mathbf{x}_{t+1}=\arg\min_{\mathbf{u}\in\mathcal{X}}\left\{\left\langle\widetilde{\mathbf{g}}_{t},\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}_{t}\right\|^{2}\right\}$ $\triangleright$ If $\mathcal{X}=\mathbb{R}^{d_{x}}$, then $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\widetilde{\mathbf{g}}_{t}$
@@
[…]
@@
The difference between Algorithm 3 and Algorithm 1, is that both the inner and outer loops sample mini-batch gradients. The inner loop guarantee is the same regardless of the inner-batch size $b_{\mathrm{in}}$, since for strongly-convex objectives it is possible to prove the same convergence rate guarantee for DP optimization in any case (as further discussed in Appendix B). As to the outer loop ($\mathbf{x}_{t+1}$ update rule), we apply standard concentration bounds to argue about the quality of the gradient estimates — hence the additive $1/b_{\mathrm{out}}$ factor — and rely on our analysis of the outer loop with inexact gradients (which are now even less exact due to sampling stochasticity). We remark that compared to the well-known analysis of Ghadimi et al. (2016) for mini-batch constrained nonconvex optimization, we derive high probability bounds without requiring several re-runs of the algorithm. We further remark that mini-batch sampling with replacement is analyzed for simplicity, though the same guarantees (up to constants) can be derived for sampling without replacement, at the cost of a more involved analysis.
@@ 6 Generalizing from ERM to population loss
Under Assumptions 2.1 and 2.2, if the preconditions of Theorem 4.1 hold, then Algorithm 1 is $(\epsilon,\delta)$-DP, and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
$\left\|\mathcal{G}_{F_{\mathcal{P}},\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha=\widetilde{\mathcal{O}}\bigg(K_{1}\Big(\frac{\sqrt{d_{x}}}{\epsilon n}\Big)^{1/2}+K_{2}\Big(\frac{\sqrt{d_{y}}}{\epsilon n}\Big)^{1/3}+\mathrm{gen}_{n}\bigg)~,$
where $K_{1}=\mathcal{O}(\Delta_{F}^{1/4}\ell^{1/4}\kappa^{5/4}),~K_{2}=\mathcal{O}(\Delta_{F}^{1/6}\ell^{5/6}\kappa^{11/6})$, and
$\mathrm{gen}_{n}=\widetilde{\mathcal{O}}\bigg(\frac{\ell\kappa(\ell\kappa+1)(\sqrt{d_{x}}+\kappa)}{\sqrt{n}}+\frac{\ell^{1/2}\kappa^{3/2}(\ell\kappa+1)}{n^{1/4}}\bigg)~.$
Similarly, if the preconditions of Theorem 5.1 hold, then for any batch sizes $b_{\mathrm{in}},b_{\mathrm{out}}\in[n]$, Algorithm 3 is $(\epsilon,\delta)$-DP, and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
$\left\|\mathcal{G}_{F_{\mathcal{P}},\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha=\widetilde{\mathcal{O}}\bigg(K_{1}\Big(\frac{\sqrt{d_{x}}}{\epsilon n}\Big)^{1/2}+K_{2}\Big(\frac{\sqrt{d_{y}}}{\epsilon n}\Big)^{1/3}+\frac{\ell\kappa}{b_{\mathrm{out}}}+\mathrm{gen}_{n}\bigg)~.$
The proof of Theorem 6.1 relies on establishing uniform convergence of the empirical hypergradients to the population hypergradients (Lemma 8.9), which as far as we know does not appear in prior literature and therefore may be of independent interest. This further implies uniform convergence of projected gradients, since the gradient mapping is non-expansive.
@@ 7 Application: private regularization hyperparameter tuning
A second approach to hyperparameter tuning is based on differentiable programming: instead of predetermining a finite set of possible values for $\omega$, we can solve Eq. (4) using BO methods, namely “differentiate through $\omega$”, and update it accordingly on the fly. This approach was pioneered by Bengio (2000), and was further developed by multiple works (Maclaurin et al., 2015; Franceschi et al., 2017, 2018; Lorraine et al., 2020; Engstrom et al., 2025). However, we are not aware of a private variant of this methodology, and our goal is to provide one here.
@@
$\omega_{t+1}$ $=\arg\min_{u\geq 0}\Big\{u\left[\nabla_{\omega}f(\omega_{t},\theta_{t}^{\lambda})+\lambda(\nabla_{\omega}g(\omega_{t},\theta_{t}^{\lambda})-\nabla_{\omega}g(\omega_{t},\theta_{t}))+\mathcal{N}(0,\sigma^{2})\right]+\tfrac{1}{2\eta}(u-\omega_{t})^{2}\Big\}$
$=\arg\min_{u\geq 0}\left\{u\left[0+\lambda(R(\theta_{t}^{\lambda})-R(\theta_{t}))+\mathcal{N}(0,\sigma^{2})\right]+\tfrac{1}{2\eta}(u-\omega_{t})^{2}\right\}$
@@ 8 Proofs
Note that the two sub-problems solved by Algorithm 1 are strongly-convex and admit Lipschitz components over $\mathcal{Y};~g(\mathbf{x},\cdot)$ by assumption, and $f+\lambda g$ by combining this with the smoothness/Lipschitzness of $f$, as follows:
@@
Although the rate in Theorem 8.2 appears in prior works such as (Bassily et al., 2014; Feldman et al., 2020), it is typically manifested through a bound in expectation (and in terms of function value) as opposed to with high probability, required for our purpose. We therefore, for the sake of completeness, provide a self-contained proof of Theorem 8.2 in Appendix B.
@@
If $\lambda\geq\max\{\frac{2L_{1}^{g}}{\mu_{g}},\frac{L_{0}^{f}}{L_{0}^{g}}\}$, then $\widetilde{\mathbf{y}}_{t}$ and $\widetilde{\mathbf{y}}_{t}^{\lambda}$ (as appear in Algorithm 1) satisfy with probability at least $1-\gamma:$
@@
where the second bound holds for $\lambda\geq\frac{L_{1}^{f}}{L_{1}^{g}}$ . ∎
@@
Noting that $\nabla^{2}_{xy}\mathcal{L}_{\lambda}\preceq 2\lambda L_{1}^{g}$ and $\nabla^{2}_{yy}\mathcal{L}_{\lambda}\succeq\lambda\mu_{g}/2$ everywhere, hence $\nabla^{2}_{yy}\mathcal{L}_{\lambda}]^{-1}\preceq 2/\lambda\mu_{g}$ we get that
@@
As $\nu_{0},\dots,\nu_{T-1}\overset{iid}{\sim}\mathcal{N}(0,\sigma^{2}I)$, a standard Gaussian norm bound (cf. Vershynin 2018, Theorem 3.1.1) ensures that with probability at least $1-\gamma$, for all $t\in\{0,1,\dots,T-1\}:~\left\|\nu_{t}\right\|^{2}\lesssim d\sigma^{2}\log(T/\gamma)\lesssim\frac{\alpha^{2}}{64}$. We therefore condition the rest of the proof on the highly probable event under which this uniform norm bound indeed holds. We continue by introducing some notation. We denote $\widetilde{\nabla}_{t}=\widetilde{\nabla}{h}(\mathbf{x}_{t})+\nu_{t}$, and $\delta_{t}:=\widetilde{\nabla}_{t}-\nabla h(\mathbf{x}_{t})$. We further denote
@@
It holds that $\langle\widetilde{\nabla}_{t},\rho_{t}\rangle\geq\left\|\rho_{t}\right\|^{2}$.
@@
$\left\langle\widetilde{\nabla}_{t}+\frac{1}{\eta}(\mathbf{x}_{t+1}-\mathbf{x}_{t}),\mathbf{u}-\mathbf{x}_{t+1}\right\rangle\geq 0~.$
@@
With the lemmas above in hand, we are now ready to prove Proposition 4.7. Note that by construction, the algorithm returns the index $t$ with minimal $\left\|\rho_{t}\right\|$. Further note that $\left\|\rho_{t}-\mathcal{G}_{t}\right\|\leq\left\|\delta_{t}\right\|$ by Lemma C.1, thus
@@
where the second inequality is due to Lemma C.1, the third due to Lemma 4.4.b, and the last by our assignment of $\lambda$. It therefore remains to bound $\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|$.
@@
$\beta=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}\right)=\widetilde{\mathcal{O}}\left(\frac{\ell^{5/2}\kappa^{11/2}\Delta_{F}^{1/2}\sqrt{d_{y}}}{\alpha^{2}\epsilon n}\right)$ (12)
@@
If $\lambda\geq\max\{\frac{2L_{1}^{g}}{\mu_{g}},\frac{L_{0}^{f}}{L_{0}^{g}},\frac{L^{f}_{1}}{L^{g}_{1}}\}$, then there is $\beta_{b}=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}+\frac{\ell\kappa}{b}\right)$ such that with probability at least $1-\gamma/2$, $\mathbf{g}_{t}^{B}:=\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};B_{t})+\lambda\left(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};B_{t})-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t};B_{t})\right)$ satisfies for all $t\in\{0,\dots,T-1\}:~\|\nabla\mathcal{L}^{*}_{\lambda}(\mathbf{x}_{t})-\mathbf{g}^{B}_{t}\|\leq\beta_{b}$.
@@
following the same proof as Lemma 4.5 in Section 8.1, by replacing Theorem 8.2 by the mini-batch Theorem B.1 (whose guarantee holds regardless of the inner batch size) .
To bound the second summand, note that $\|\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};\xi)+\lambda(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};\xi)-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t};\xi))\|\leq M=\mathcal{O}(\ell\kappa)$ for every $\xi\in\Xi$, by Lemma 4.6. Hence, $\mathbf{g}^{B}_{t}$ is the average of $b$ independent vectors bounded by $M$, all with the same mean, and therefore a standard concentration bound (cf. Jin et al. 2019) ensures that $\|\mathbf{g}_{t}^{B}-\mathbb{E}[\mathbf{g}_{t}^{B}]\|=\widetilde{\mathcal{O}}(M/b)$ with probability at least $1-\gamma/4$, which completes the proof. ∎
@@
[…]
@@
where the second inequality is due to Lemma C.1, the third due to Lemma 4.4, and the last by our assignment of $\lambda$. It therefore remains to bound $\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|$.
@@
$\beta_{b}=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}+\frac{\ell\kappa}{b}\right)=\widetilde{\mathcal{O}}\left(\frac{\ell^{5/2}\kappa^{11/2}\Delta_{F}^{1/2}\sqrt{d_{y}}}{\alpha^{2}\epsilon n}+\frac{\ell\kappa}{b}\right)~,$ (16)
@@
The proof is based on the following lemma that establishes uniform convergence of hypergradients.